# Multiwavelength Observations of the Apparently Non-repeating FRB 20250316A

### v1

@@ 3.3.1 WFST
The WFST images were reduced using the standard optical reduction procedures. After extracting objects using SExtractor 111https://sextractor.readthedocs.io/en/latest/Introduction.html, we cross-matched them with Gaia and refined the WCS of the images using SCAMP222https://www.astromatic.net/software/scamp/. Photometric zero points were then estimated with Pan-STARRS333https://outerspace.stsci.edu/display/PANSTARRS -band magnitudes. In order to examine the details of the host, we stacked the images using SWarp 444https://www.astromatic.net/software/swarp/ with the median values. The stacked image is presented in Figure 3. The positions of FRB 20250316A (red), EP J120944.2+585060 (yellow), and CXOU J120943.7+585051 (cyan) are overplotted. This shows that FRB 20250316A and CXOU J120943.7+585051 are located in two extended regions of the host galaxy, possibly star-forming regions. The center of the extended region near FRB 20250316A is offset by 1.0 arcsec from the FRB localization. The corresponding projected physical distance is pc, at a distance of 40 Mpc.
After this, we searched for possible optical counterparts of FRB 20250316A in individual images and the stacked image, using the DESI/Legacy Survey (DESI/LS)555https://www.legacysurvey.org/ band image as a template. For individual images, we first registered them to the stacked DESI/LS image using reproject in Python. The images were then subtracted with HOTPANTS 666https://github.com/acbecker/hotpants with a kernel size of 1.5 times the WFST FWHM. There were no significant objects at the position of FRB 20250316A in individual subtracted images. The upper limits are estimated as , where ZP is the zero point of each image and is the RMS of the FRB 20250316A field, with a radius of 90 arcsec, in the residual image. The 5 upper limits are listed in Table 3 and Table LABEL:tb:opt1img. An upper limit from the stacked image was estimated with a similar process. Limited by the seeing and the depth of the DESI/LS image, the subtracted image has a limiting magnitude of mag around the FRB 20250316A field.
@@ 3.3.2 SVOM/VT
We first stacked the images obtained on 2025-03-23 and 2025-03-25 individually. For image stacking, we followed the standard procedure using IRAF v2.16 (D. Tody, 1986, 1993): (1) Images with abnormally high background levels or distorted stellar profiles were excluded. The remaining images underwent bias, dark, and flat-field corrections. (2) Precise astrometric alignment was performed with astrometry.net (D. Lang et al., 2010; P. Joseph & A. Dagore, 2024) for each image relative to a chosen reference frame, based on the positions of common bright stars. (3) The aligned images were then stacked using the IRAF.IMCOMBINE task with an average combination method, incorporating sigma-clipping to effectively reject cosmic rays and bad pixels, thereby enhancing the signal-to-noise ratio of the final stacked images. For both the VTB and VTR bands of SVOM/VT, we stacked 101 and 97 frames obtained on 2025-03-23 and 2025-03-25, respectively, with each individual frame having an exposure time of 60 seconds (see Table 3 and Table LABEL:tb:opt1img). We then searched for possible optical counterparts with a similar process to that of WFST. Because SVOM/VT observations were taken on two different days, and used custom-designed filters ( nm for VTB and nm for VTR), we utilized the images from SVOM/VT itself as templates to perform image subtraction. Since signals temporally closer to the burst are expected to be more significant, the last single images and the stacked images from 2025-03-25 were used as templates, for single-image and stacked-image subtraction, respectively. The results are summarized in Table 3 and Table LABEL:tb:opt1img.
@@ 4.1 One-off or Repeating
One of the most critical questions in FRB research is whether all FRBs repeat. As a nearby FRB, FRB 20250316A provides a valuable opportunity to examine this issue. Observationally, it was detected by CHIME after 269.77 hours of exposures over six years (FRB Collaboration et al., 2025a), corresponding to an FRB rate of 0.004 per hour with a 1 range of per hour ( erg Hz-1). Following its discovery, FAST conducted 13.08 hours of follow-up observations over approximately one week (Section 3.1). Assuming a Poisson distribution, this indicates a 3 upper limit of 6.6/13.08=0.5 per hour ( erg Hz-1). The left panel of Figure 4 presents FRB rates estimated with CHIME (red dot) and the upper limit from FAST (red triangle). For comparison, we overplot the energy functions of repeating FRBs. The energy functions of repeating FRBs are empirically fitted as a power law, with a slope of , spanning a range from to (C. J. Law et al., 2017; K. Gourdji et al., 2019). A flatter high-energy component has sometimes been proposed for repeating FRBs, such as FRB 20121102A (D. Li et al., 2021; D. M. Hewitt et al., 2022), FRB 20200120E (S. B. Zhang et al., 2024), FRB 20201124A (Y.-K. Zhang et al., 2022; F. Kirsten et al., 2024), and FRB 20220912A (O. S. Ould-Boukattine et al., 2024). The flatter component typically extends over orders of magnitude in spectral energy. A comparison between FRB 20250316A and repeating FRBs reveals two possible scenarios. First, the typical power law with a slope of (red solid line) predicts an FRB rate five orders of magnitude larger than the FAST 3 upper limit. Second, scaling the energy functions of repeating FRBs to the rate of FRB 20250316A at its spectral energy (blue lines) shows that repeating FRBs predict rates 2–8 orders of magnitude higher than the FAST upper limit.
@@
On the other hand, some repeating FRBs have 10 bursts, with FRB rates of bursts per hour at a completeness of Jy ms (Chime/Frb Collaboration et al., 2023). Their energy functions cannot be examined in detail. In order to compare with them, we estimated the FRB rate of FRB 20250316A by combining observations from multiple telescopes. A cumulative FRB rate with multiple telescopes can be estimated as
@@
where is the energy function of FRBs, iterates over different telescopes, and is the threshold of each telescope. To compare with the CHIME repeating FRBs, we adopt following Chime/Frb Collaboration et al. (2023). Taking FAST, CHIME and HyperFlash observations into account and assuming a slope of , we obtain a combined FRB 20250316A rate of hour-1 (red region in the right panel of Figure 4). This region lies below all CHIME repeating FRBs (blue histogram). Using the flattest slope of repeating FRBs, (from the high energy component of FRB 20200120E), the rate is estimated to be hour-1 (gray region), still below all CHIME repeating FRBs.
@@
We caution, however, that this process assumes the energy function of repeating FRBs is stable, whereas temporal variations are possible. In addition, we use only the “gold” sample from Chime/Frb Collaboration et al. (2023). The “silver” sample, which may be contaminated by chance coincidence, may yield CHIME repeating FRBs with a lower rate. In summary, FRB 20250316A is different from the currently known repeating FRB sample and is highly likely to be a one-off event.
@@ 4.2 Constraints on multiwavelength counterparts
Persistent X-ray emission is also expected in various models. In the magnetar scenario, outbursts of persistent X-ray emission are usually observed during periods of enhanced bursting and flaring activity (V. M. Kaspi & A. M. Beloborodov, 2017; F. Coti Zelati et al., 2018; N. Rea & D. De Grandis, 2025), as in the case of the X-ray burst from the Galactic magnetar SGR 1935+2154 associated with FRB 20200428D (C. D. Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020). Typical long-term X-ray outbursts of magnetars might last several months to years, reaching a maximum luminosity of erg s-1 (most of the luminosity is expected to be released in neutrinos (J. A. Pons & N. Rea, 2012)). Given our upper limit of erg s-1, we cannot place meaningful constraints on such persistent magnetar outbursts. However, if FRBs are produced by a very young magnetar with fast rotation, one might expect a large amount of rotational power () that could support an X-ray counterpart in the form of pulsed synchrotron emission or a wind nebula. For example, by assuming: 1) a magnetic field of at least G (required for the magnetar to show large X-bursts related to the FRB radio counterpart), and 2) an X-ray efficiency of (in line with rotationally powered young pulsars), our deep upper limit implies a maximum rotational power for the young magnetar of erg s-1 (assuming a NS mass of 1.4 and radius of 12 km). Under these assumptions, the putative young magnetar responsible for FRB 20250316A would have a spin period larger than P and a characteristic age longer than 1 yr, possibly in line with the magnetar population of this star-forming galaxy.
Recent observations of FRB environments and possible periodicities imply that FRBs may relate to binary systems (K. Ioka & B. Zhang, 2020; F. Y. Wang et al., 2022; R. Anna-Thomas et al., 2023; Y. Li et al., 2025). Thus, neutron star binary systems, especially high-mass X-ray binaries (HMXBs) and ultraluminous X-ray sources (ULXs), might also be related to FRBs.
In Figure 5, we compare the persistent X-ray luminosity upper limit of FRB 20250316A with those expected from different neutron star classes, both in isolation and in binary systems. The Chandra X-ray upper limit of FRB 20250316A is fainter than nearly all ULXs, and is compatible with an X-ray binary or a magnetar. This disfavors a direct connection between FRBs and ULXs, although the strong variability observed in some ULXs prevents us from completely ruling out these systems as possible progenitors of FRB 20250316A.
@@
Ultra-relativistic outflow: If the deceleration time of the fireball occurs prior to the first WXT observation (i.e., for ), the first WXT point at day provides the most stringent constraint on the kinetic energy, yielding erg. In this case, the first FXT point at day further constrains the jet opening angle to , requiring a jet break time earlier than day to keep the X-ray flux below the FXT measurements. The predicted X-ray lightcurve is presented as a blue solid line. The corresponding optical emission (orange solid line) is consistent with the upper limits from SVOM and WFST, as well as the absence of radio counterparts (T. An et al., 2025).
@@
Low Lorentz factor outflow: If the fireball decelerates around day (near the FXT points), the inferred Lorentz factor is relatively low, . In this scenario, the FXT upper limits do not constrain the jet kinetic energy. Instead, the optical upper limits from SVOM and WFST dominate, requiring erg (orange dotted line). The corresponding X-ray lightcurve is shown as the blue dotted line.
@@
4.3 Properties of CXOU J120943.7+585051
Although the precise location of CXOU J120943.7+585051/EP J120944.2+585060 provided by Chandra rules out its association with FRB 20250316A, the nature of the X-ray source remains of interest. The source was detected by FXT in observations spanning over 50 days. The keV light curve extracted at the Chandra position shows moderate variability (see Table 2 and Figure 6). However, a test against a constant flux model (p=0.419) revealed no statistically significant variability. The overall flux remained stable within uncertainties, as also indicated by the late observation in May.
With an estimated flux of inferred from the Chandra observation, this source, if associated with the host galaxy at 40 Mpc, corresponds to a luminosity of . This places it within the typical luminosity range for ultraluminous X-ray Sources (ULXs), which are thought to be powered by accretion onto a compact object (either a stellar-mass black hole, neutron star, or potentially an intermediate-mass black hole) at a luminosity comparable to or above the Eddington luminosity of stellar black holes (see P. Kaaret et al. 2017 for a review). Alternatively, if CXOU J120943.7+585051 is not associated with the FRB’s host galaxy, it could be a foreground Galactic source (such as a cataclysmic variable) or a background extragalactic object like an active galactic nucleus (AGN). Distinguishing between these possibilities would require multiwavelength observations.
4.4 Probability of Chance Coincidence
The X-ray source CXOU J120943.7+585051 was detected by Chandra within the FXT error circle. Given its relatively large separation of 7 arcseconds from the FRB position, the source is unlikely to be physically associated with FRB 20250316A. To quantify the likelihood of such a chance alignment, we calculated the probability of a random, unrelated ULX appearing within 10 arcseconds of the FRB position in a galaxy. It should be noted that alternative origins for CXOU J120943.7+585051, such as a foreground Galactic X-ray source or a background Active Galactic Nucleus (AGN), cannot be fully ruled out.
A 10-arcsecond angular separation at 40 Mpc corresponds to a projected physical radius of 1.94 kpc. The likelihood of finding an unrelated ULX within this region depends critically on the host galaxy properties and the surface density of ULXs. For the general galaxy population, the average number of ULXs ( ) is reported to be (K. Kovlakas et al., 2020). For a typical galactic disk of radius 15 kpc, the expected number of ULXs in a 1.94 kpc radius circle is . Using a Poisson distribution, this corresponds to a chance probability of 1.6%. However, this probability is significantly enhanced in star-forming regions or star-forming galaxies, where ULXs are substantially more prevalent (D. A. Swartz et al., 2011; H. Feng & R. Soria, 2011). In such active environments, the density of ULXs could be higher by a factor of 5 to 10, corresponding to a chance probability of approximately –% to detect an unrelated ULX-like source within 10 arcseconds in a star-forming galaxy. Such a non-negligible probability highlights the importance of precise localization, even when an apparently related counterpart candidate is initially identified.
Future investigations of FRB counterparts should be cautious about contamination from unrelated sources such as LMXBs, HMXBs, ULXs, low-luminosity supermassive black holes (SMBHs), or intermediate-mass black holes (IMBHs). As discussed above, the average ULX density is about one per galaxy, with higher densities in spiral galaxies and lower densities (0.5 per galaxy) in ellipticals and dwarfs. Since a significant fraction of FRB host galaxies are star-forming spiral galaxies, the probability of having an ULX ( erg s-1) by chance coincidence is relatively high. Furthermore, given that around 20% FRBs reside in host galaxies with AGN signals (T. Eftekhari et al., 2023; K. Sharma et al., 2024), the chance coincidence probability of contamination from the SMBHs is then non-negligible. These considerations highlight the importance of arcsecond-level localization for both FRBs and their potential X-ray counterparts.

### v2

@@ 3.3.1 WFST
WFST images were reduced with standard optical procedures. After extracting objects using SExtractor (E. Bertin & S. Arnouts, 1996), we cross-matched them with Gaia and refined the WCS using SCAMP (E. Bertin, 2006). Photometric zero points were calibrated with Pan-STARRS -band magnitudes (K. C. Chambers et al., 2016). To examine host details, we stacked images using SWarp (E. Bertin et al., 2002) with median values. The stacked image is shown in the lower right panel of Figure 2, with the positions of FRB 20250316A (red), EP J120944.2+585060 (yellow), and CXOU J120943.7+585051 (cyan) overplotted. FRB 20250316A and CXOU J120943.7+585051 lie in two extended regions, likely star-forming regions of the host galaxy. The center of the star-forming region hosting FRB 20250316A was determined at (RA, Dec) = (182.434782∘, 58.848778∘) using the centroid_com function from Photutils. The FRB lies 1.1′′ from this centroid (215 pc at 40 Mpc), consistent with CHIME/FRB Collaboration et al. (2025) and P. K. Blanchard et al. (2025b). To assess systematic uncertainties, we also measured the offset between the FRB and the peak of the star-forming region, yielding 0.8′′ (157 pc).
After this, we searched for possible transient optical counterparts of FRB 20250316A in individual and the stacked image, using the DESI/Legacy Survey (DESI/LS)111https://www.legacysurvey.org/ band image as a template. Individual images were aligned to the DESI/LS stack with reproject (T. Robitaille et al., 2020) in Python, then subtracted using HOTPANTS (A. Becker, 2015) with a kernel size of 1.5 times the WFST FWHM. No significant objects were found at the FRB 20250316A position in individual subtracted images. The upper limits were estimated as , where ZP is the zero point of each image and is the RMS within 90′′ of the FRB in the residual image. The 5 upper limits are listed in Table 3 and Table LABEL:tb:opt1img. An upper limit of mag was estimated from the stacked image using a similar process, limited by the seeing and DESI/LS depth. This is generally consistent with the limits of FRB optical counterpart from MMT telescope ( on 2025 March 23) and Gemini ( on 2025 March 24) (CHIME/FRB Collaboration et al., 2025).
@@ 3.3.2 SVOM/VT
We first stacked the images obtained on 2025-03-23 and 2025-03-25 individually, following the standard procedure using IRAF v2.16 (D. Tody, 1986, 1993): (1) Images with abnormally high backgrounds or distorted stellar profiles were excluded. The remaining images underwent bias, dark, and flat-field corrections. (2) Precise astrometric alignment was performed with astrometry.net (D. Lang et al., 2010; P. Joseph & A. Dagore, 2024) relative to a chosen reference frame using common bright stars. (3) The aligned images were then stacked with IRAF.IMCOMBINE task with an average combination method, incorporating sigma-clipping to reject cosmic rays and bad pixels, improving the signal-to-noise ratio of the final stacked images. For both the VTB and VTR bands, we stacked 101 and 97 frames from 2025-03-23 and 2025-03-25, respectively, each with 60 s exposure. We then searched for possible optical counterparts with a similar process to that of WFST. Because SVOM/VT observations were taken on two different days, and used custom-designed filters ( nm for VTB and nm for VTR), we utilized the images from SVOM/VT itself as templates to perform image subtraction. Since signals temporally closer to the burst are expected to be more significant, the last single and the stacked images from 2025-03-25 were used as templates, for single and stacked image subtraction, respectively. The upper limits are summarized in Table 3 and Table LABEL:tb:opt1img.
@@ 4 Physical Implications and Discussions
4.1 Probability of Chance Coincidence
The X-ray source CXOU J120943.7+585051 was detected by Chandra within the FXT error circle. Given its relatively large separation of 7 arcseconds from the FRB position, the source is unlikely to be physically associated with FRB 20250316A. To quantify the likelihood of such a chance alignment, we calculated the probability of chance coincidence in this section. It should be noted that alternative origins for CXOU J120943.7+585051, such as a foreground Galactic X-ray source or a background AGN, cannot be fully ruled out.
Firstly, we estimate the probability to have an X-ray source within a circle with a 10′′ radius on average. Based on the X-ray source density (A. Masini et al., 2020) at EP/FXT limiting flux, in keV, the averaged number of X-ray sources within the EP/FXT angular resolution is .
Secondly, the density of X-ray sources within galaxies may be enhanced. Thus, we also calculated the probability of a random, unrelated ULX appearing within 10 arcseconds of the FRB position in a galaxy. A 10-arcsecond angular separation at 40 Mpc corresponds to a projected physical radius of 1.94 kpc. The likelihood of finding an unrelated ULX within this region depends critically on the host galaxy properties and the surface density of ULXs. For the general galaxy population, the average number of ULXs ( ) is reported to be (K. Kovlakas et al., 2020). For a typical galactic disk of radius 15 kpc, the expected number of ULXs in a 1.94 kpc radius circle is , giving a 1.6% chance probability using a Poisson statistics. However, this probability is significantly enhanced in star-forming regions or star-forming galaxies, where ULXs are substantially more prevalent (D. A. Swartz et al., 2011; H. Feng & R. Soria, 2011). In such active environments, the density of ULXs could be higher by a factor of 5 to 10, corresponding to a chance probability of approximately –% to detect an unrelated ULX-like source within 10 arcseconds in a star-forming galaxy.
In addition, for even nearby FRBs, low-mass X-ray binaries (LMXBs) and high-mass X-ray binaries (HMXBs) may contribute to contamination. Using a similar method as above, and including the luminosity functions of LMXBs, HMXBs, and ULXs (S. Wang et al., 2016; M. Gilfanov et al., 2022), we estimate the average number of unrelated LMXBs, HMXBs, and ULXs expected within the angular resolution of EP/FXT (10′′) and Chandra (1′′) for different luminosity distances , as shown in Figure 4. It shows that the chance of contamination from LMXBs, HMXBs, and ULXs is about 1% for EP/FXT and 0.01% for Chandra. Furthermore, given that around 20% FRBs reside in host galaxies with AGN signals (T. Eftekhari et al., 2023; K. Sharma et al., 2024), the contamination from AGNs or intermediate-mass black holes (IMBHs) is then non-negligible, especially for higher redshifts. These considerations highlight the importance of arcsecond-level localization for both FRBs and their potential X-ray counterparts, although EP/FXT candidates remain useful in most cases.
@@ 4.2 One-off or Repeating
One of the most critical questions in FRB research is whether all FRBs repeat. As a nearby FRB, FRB 20250316A provides a valuable opportunity to examine this issue. Observationally, it was detected by CHIME after 269.77 hours of exposures over six years (CHIME/FRB Collaboration et al., 2025), corresponding to an FRB Poisson rate of 0.004 per hour with a 1 range of per hour ( Jy ms, corresponding to erg Hz-1). FAST conducted 13.08 hours of follow-up observations over approximately one week (Section 3.1). Assuming a Poisson distribution, this indicates a 3 upper limit of ( mJy ms, corresponding to erg Hz-1). The left panel of Figure 5 presents FRB rates estimated with CHIME (red dot) and the upper limit from FAST (red triangle). For comparison, we overplot the energy functions of repeating FRBs. The energy functions of repeating FRBs are empirically fitted as a power law, with a slope of , spanning a range from to (C. J. Law et al., 2017; K. Gourdji et al., 2019). A flatter high-energy component has sometimes been proposed for repeating FRBs, such as FRB 20121102A (D. Li et al., 2021; D. M. Hewitt et al., 2022), FRB 20200120E (S. B. Zhang et al., 2024), FRB 20201124A (Y.-K. Zhang et al., 2022; F. Kirsten et al., 2024), and FRB 20220912A (O. S. Ould-Boukattine et al., 2024). Their energy functions are shown as blue lines, with the lower index, higher index, and the break spectral energy labeled. The flatter component typically extends over orders of magnitude in spectral energy. A comparison between FRB 20250316A and repeating FRBs reveals two possible scenarios. First, the typical power law with a slope of (red solid line) predicts an FRB rate of above the FAST limiting fluence of , more than five orders of magnitude larger than the FAST 3 upper limit. Second, scaling the energy functions of repeating FRBs to the rate of FRB 20250316A at its spectral energy (blue lines) shows that repeating FRBs predict rates 3–7 orders of magnitude higher than the FAST upper limit.
We note that repeating FRBs are usually clustered, with waiting times often described by Weibull distributions (N. Oppermann et al., 2018; L. C. Oostrum et al., 2020; Chime/Frb Collaboration et al., 2020; K. M. Rajwade et al., 2020; F. Kirsten et al., 2024). The Weibull shape parameter characterizes clustering: bursts are clustered if , with typical values – (N. Oppermann et al., 2018; L. C. Oostrum et al., 2020; R. Luo et al., 2020). Using the simulation method of R. Luo et al. 2020, we obtained FRB rate upper limits of 2.0, 0.7, and 0.5 (at FAST limiting fluence) for , 0.5, and 0.9, respectively—still 2–7 orders of magnitude below predictions from repeating FRB energy functions.
@@
On the other hand, some repeating FRBs have 10 bursts, with FRB rates of bursts per hour at a completeness of Jy ms (Chime/Frb Collaboration et al., 2023b). Their energy functions cannot be examined in detail. In order to compare with them, we estimated the FRB rate of FRB 20250316A by combining observations from multiple telescopes. A cumulative FRB rate with multiple telescopes can be estimated as
@@
where is the energy function of FRBs, iterates over different telescopes, and is the threshold of each telescope. To compare with the CHIME repeating FRBs, we adopt following Chime/Frb Collaboration et al. (2023b). Taking into account the observations of FAST, CHIME, and HyperFlash—a 25–32 m European radio telescope network that observed FRB 20250316A for 120 hours at L-band with a limiting fluence of 15 Jy ms (O. S. Ould-Boukattine et al., 2025; CHIME/FRB Collaboration et al., 2025)—and assuming a slope of , we obtain a combined FRB 20250316A rate of hour-1 (red region in the right panel of Figure 5). This region lies below all CHIME repeating FRBs (blue histogram). Using the flattest slope of well-studied repeating FRBs, (FRB 20200120E, S. B. Zhang et al. 2024), the rate is estimated to be hour-1 (gray region), still below all CHIME repeating FRBs. Thus, FRB 20250316A is different from the currently known repeating FRB sample.
@@
CHIME/FRB Collaboration et al. 2025 explored the repeatility of this burst by a simulation, and concludes that the probability of detecting only one burst like FRB 20250316A in repeating FRBs is more significant than 5. With different methods, we have the same conclusion that FRB 20250316A is highly likely to be a one-off event. We caution, however, that this process assumes the energy function of repeating FRBs is stable, whereas temporal variations are possible (D. Li et al., 2021; S. B. Zhang et al., 2024). In addition, we use only the “gold” sample from Chime/Frb Collaboration et al. (2023b). The “silver” sample, which may be contaminated by chance coincidence, may yield CHIME repeating FRBs with a lower rate. As a result, it is hardly to totally exclude the possibility that FRB 20250316A would be detected in some distant future.
@@ 4.3 Constraints on multiwavelength counterparts
Persistent X-ray emission is also expected in various models. In the magnetar scenario, outbursts of persistent X-ray emission are usually observed during periods of enhanced bursting and flaring activity (V. M. Kaspi & A. M. Beloborodov, 2017; F. Coti Zelati et al., 2018; N. Rea & D. De Grandis, 2025), as in the case of the X-ray burst from the Galactic magnetar SGR 1935+2154 associated with FRB 20200428D (C. D. Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020). Typical long-term X-ray outbursts of magnetars might last several months to years, reaching a maximum luminosity of erg s-1 (V. M. Kaspi & A. M. Beloborodov, 2017), where most of the luminosity is expected to be released in neutrinos (J. A. Pons & N. Rea, 2012). Given our upper limit of erg s-1, we cannot place meaningful constraints on such persistent magnetar outbursts. However, if FRBs are produced by a very young magnetar with fast rotation, one might expect a large amount of rotational power () that could support an X-ray counterpart in the form of pulsed synchrotron emission or a wind nebula. For example, by assuming: 1) a magnetic field of at least G (required for the magnetar to show large X-bursts related to the FRB radio counterpart, R. Turolla et al. 2015), and 2) an X-ray efficiency of (in line with rotationally powered young pulsars, A. Possenti et al. 2002), our deep upper limit implies a maximum rotational power for the young magnetar of erg s-1 (assuming a NS mass of 1.4 and radius of 12 km). Under these assumptions, the putative young magnetar responsible for FRB 20250316A would have a spin period larger than P and a characteristic age longer than 1 yr, possibly in line with the magnetar population of this star-forming galaxy.
Recent FRB environment studies and possible periodicities imply a possible link to binary systems (K. Ioka & B. Zhang, 2020; F. Y. Wang et al., 2022; R. Anna-Thomas et al., 2023; Y. Li et al., 2025; B. Zhang & R.-C. Hu, 2025). Thus, neutron star binary systems, especially LMXB, HMXBs and ULXs, might also be related to FRBs.
In Figure 6, we compare the persistent X-ray luminosity upper limit of FRB 20250316A with those expected from different neutron star classes, both in isolation and in binary systems. We also overplot upper limits from other deep X-ray observations, including the repeating FRB 20200120E (gray open arrow; A. B. Pearlman et al. 2025), as well as the non-repeating FRBs 20200430A and 20190608B (gray filled arrows; T. Eftekhari et al. 2023). The Chandra X-ray upper limit of FRB 20250316A is fainter than nearly all ULXs, while still compatible with an X-ray binary or a magnetar. It is deeper than previously reported limits for other non-repeating FRBs, but less constraining than the limit for the repeating FRB 20200120E. Although the strong variability observed in some ULXs prevents us from completely ruling out these systems as possible progenitors of FRB 20250316A, this result disfavors a direct connection between non-repeating FRBs and ULXs, in agreement with previous findings for the repeating FRB 20200120E.
@@
Ultra-relativistic outflow: If the fireball decelerates before the first WXT observation (i.e., for ), the WXT point at day provides the most stringent kinetic energy constrain, erg. The first FXT point at day then limits the jet opening angle to , requiring a jet break before day to keep the X-ray flux below the FXT measurements (see blue solid line). The corresponding optical emission (orange solid line) is also consistent with the upper limits from SVOM and WFST, as well as the absence of radio counterparts (T. An et al., 2025).
@@
Low Lorentz factor outflow: If the fireball decelerates around day (near the FXT points), the inferred Lorentz factor is relatively low, . In this scenario, the FXT upper limits do not constrain the jet kinetic energy. Instead, the optical upper limits from SVOM and WFST dominate, requiring erg. The corresponding X-ray lightcurve is shown as the blue dotted line. If we take the MMT upper limit () and and Gemini ( on 2025 March 24) into account, it requires .